what is the solution?
6x - 7y=-56
4x + 52=y

Answer by mananth(16946)

6 x -7 y = -56 .............1
4 x -1 y = -52 .............2
Eliminate y
multiply (1)by -1
Multiply (2) by 7
-6 x 7 y = 56
28 x + -7 y = -364
Add the two equations
22 x = -308.00
/ 22
x = -14
plug value of x in (1)
6 x + -7 y = -56
-84 + -7 y = -56
-7 y = -56 + 84
-7 y = 28
y = -4